OpenStudy (anonymous):

A rectangular prism has a volume of 3m^3, a length of 30cm, and a width of 40cm. What is the height of the prism?

OpenStudy (superdavesuper):

its a rectangular prism....wont its volume = length x width x height?

OpenStudy (anonymous):

The volume of a rectangular prism can be found by the formula: volume=length*width*height

OpenStudy (anonymous):

3m^3 = 30cm * 40cm* ht.cm

OpenStudy (superdavesuper):

check ur units....again @pmcafee ;)

OpenStudy (anonymous):

3m^3 * 1,000,000/1m^3 = 3,000,000cm^3 3,000,000cm^3 = 1200cm^2(ht) 3,000,000cm^3/1200cm^2 = 2500cm (ht.) ???

OpenStudy (superdavesuper):

yup yup....would have less 00000000000 if u go w m but thats the right ans :)

OpenStudy (anonymous):

One of the answers was 25m. So, I converted 2500cm * 1m/100cm = 25m

OpenStudy (superdavesuper):

right on :)